Case Study : Nonprofit Hospital Accounting Equation The expanded accounting equation for a nonprofit is as follows: Assets = Liabilities + Net assets + Revenues – Expenses Assignment: a) Indicate all financial effects of each transaction on the expanded accounting equation See transaction 1 below- effect on equation is increase of Assets and Liabilities for $5M each - everything is in excel, below it shows transaction descriptions of 1-10, then Assets, Liabilities, Net Assets, Revenues, Expenses b) After recording the transactions, total the columns and prepare an abbreviated balance sheet and income statement below c) What is net income? d) What is Total Net Assets?

1. Hospital borrowed $5 million from bank. - excel shows $5M in assets and $5M in liabilities

2. Hospital purchased $2 million of supplies on account.

3. 50% of the supplies in transaction #2 were subsequently used for patient care.

4. Hospital billed patients $10 million for services provided.

5. Hospital paid employees $3 million wages. 6. Hospital collected $7 million for services billed in #4.

7. Hospital paid $2 million to suppliers for supplies purchase in #2.

8. Hospital paid bank back $1 million of loan plus $100,000 of interest.

9. Hospital paid utility bills of $1 million.

10. Hospital recorded $3 million of depreciation on equipment. Totals
Answer #1

Answer A:

Expanded Accounting Equation for non-profit is as under:
Assets (A) = Liabilities (L) + Net Assets (NA) + Revenue (R) - Expenses (E)
A L NA R E
Transaction 1 5 5
Transaction 2 2 2
Transaction 3 1 -1
Transaction 4 10 10
Transaction 5 -3 3
Transaction 6 0
Transaction 7 -2 -2
Transaction 8 -1.1 -1 0.1
Transaction 9 -1 1
Transaction 10 -3 3
5.9 4 0 10 8.1

Answer B:

INCOME STATEMENT (Amt in $ Million)
Dr. Cr.
PARTICULARS AMOUNT PARTICULARS AMOUNT
Supplies Purchase 2 Receipts from Patients 10
Wages 3 Supplies in Stock 1
Interest on Bank Loan 0.1
Utility Bills Expense 1
Depreciation Expense 3
Net Income 1.9
11 11
ABBREVIATED BALANCE SHEET (Amt in $ Million)
LIABILITIES AMOUNT ASSETS AMOUNT
Bank Loan 4 Accumulated Depreciation -3
Supplies Payable 0 Receipts Receivable 3
Net Income 1.9 Supplies in Stock 1
Bank Balance 4.9
5.9 5.9

Answer C: As seen in Income Statement, Net Income of the non-profit hospital is $ 1.9 Million.

Answer D:

Expanded Accounting Equation for non-profit is as under:

Assets (A) = Liabilities (L) + Net Assets (NA) + Revenue (R) - Expenses (E)

Thus, NA = A - L - R + E,

NA = 5.9 - 4 - 10 + 8.1 = 0

Thus, Net Assets of the Non-Profit Hospital is "0".